OpenStudy (anonymous):

Use the Double-Angle Formulas to rewrite cos4x in terms of cos x HELP. I knw the steps but ill explain more in comments...

OpenStudy (anonymous):

I get to 2(cos2x)^2-1...

OpenStudy (anonymous):

@primeralph :/ please

OpenStudy (primeralph):

This is the kind of thing where you just keep going until you're done. Not hard.

OpenStudy (anonymous):

but i dont see how you do the next step

OpenStudy (anonymous):

the next thing i should get is...2(2cos2x - 1)^2 - 1

OpenStudy (anonymous):

Recall that \[\cos(A+B) = \cos(A)\cos(B) -\sin(A)\sin(B)\] so cos(4x) = cos (2x + 2x) right? now substitute. and solve

OpenStudy (primeralph):

Do the same for cos2x.

OpenStudy (anonymous):

how is my question

OpenStudy (primeralph):

Tell me what you get.

OpenStudy (anonymous):

ok...thinking

OpenStudy (anonymous):

2(cos^2u-1)-1

OpenStudy (anonymous):

oh parenthhesis squared

OpenStudy (primeralph):

u?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

sorry....

OpenStudy (primeralph):

Then expand or something.

OpenStudy (anonymous):

so then 2(4cos^4x-4cos^2x+1)-1

OpenStudy (anonymous):

distribute...

OpenStudy (anonymous):

8cos^4x-8cos^2+1?

OpenStudy (anonymous):

that was the answer :D

OpenStudy (anonymous):

i wish i could figure this out myself...thanks guys

OpenStudy (primeralph):

Hold on.

OpenStudy (anonymous):

ok

OpenStudy (primeralph):

Let me see all the options.

OpenStudy (anonymous):

that was the only option..it wasnt for a test..i dont cheat like that...this was me trying to understand an example..

OpenStudy (anonymous):

that was the answer provided

OpenStudy (primeralph):

Okay.

OpenStudy (anonymous):

thank you :)

OpenStudy (primeralph):

You're welcome.